OpenStudy (anonymous):

15b^8c^9,5b^5c^8,21b^4c^2 Find the LCM. I am having a hard time with these

OpenStudy (anonymous):

I'll try to simplify this for you...so that u can do it yourself....

OpenStudy (anonymous):

first write each as a product of the number and different variables 15b^8c^9 = 15 * b^8 * c^9 5b^5c^8 = 5 * b^5 * c^8 21b^4c^2 = 21 * b^4 * c^2

OpenStudy (anonymous):

Then find the LCM of the numbers (here they r 15, 5 and 21) ..... I hope u know how to do this LCM of these numbers is 105 Now for each variable take the one with highest power/index for b it is b^8 for c it is c^9 Now combine all these to get the LCM as 105b^8c^9